Domain and Range-

16. f(x) = 11x + 15
g(x) = -11x3 - 15x2 + 13x - 13
find g(f(6))
A. -4595
B. -52466
C. -6288203
D. -5943226
E. 639
F. 538

17. f(x) = 15x3 + 13x2 - 15x + 15
g(x) = 15x + 11
find g(f(9))
A. 639
B. -5943226
C. -52466
D. -6288203
E. 178031
F. -4595

Answer by ikleyn(52790)

16. f(x) = 11x + 15
g(x) = -11x3 - 15x2 + 13x - 13
find g(f(6))
A. -4595
B. -52466
C. -6288203
D. -5943226
E. 639
F. 538
~~~~~~~~~~~~~~~~~~

First calculate f(6) = 11*6 + 15 = 66 + 15 = 81.

Then substitute this value as x into g(x):

g(81) = -11*81^3 - 15*81^2 + 13*81 - 13 = . . .

Could you please calculate it yourself?


With the second composition the procedure is the same.